prove that Q[x]/<x-2> is isomorphic to Q as fields.
1
Expert's answer
2020-10-11T17:54:07-0400

Polynomial remainder theorem implies that each element of "\mathbb Q[x]/\langle x-2\rangle" is of the form "[f(x)]=f(x)+\langle x-2\rangle=f(x)+(x-2)\mathbb Q[x]", where "\deg f(x)<\deg(x-2)=1", and therefore, "f(x)\in\mathbb Q". Consequently, "\mathbb Q[x]/\langle x-2\rangle=\{[a]\ : \ a\in\mathbb Q\}".


Consider the map "\psi: \mathbb Q\to \mathbb Q[x]/\langle x-2\rangle, \psi(a)=[a]." Taking into account that "\psi(ab)=[ab]=[a][b]=\psi(a)\psi(b)" and "\psi(a+b)=[a+b]=[a]+[b]=\psi(a)+\psi(b)", we conclude that "\psi" is a field homomorphism. If "a\ne b", then "\psi(a)=[a]\ne [b]=\psi(b)", and consequently "\psi" is injective. Since "\psi(a)=[a]" for any "[a]\in \mathbb Q[x]/\langle x-2\rangle", "\psi" is surjective. Therefore, "\psi" is a field isomorphism. Thus the fields "\mathbb Q" and "\mathbb Q[x]/\langle x-2\rangle" are isomorphic.